OpenStudy (anonymous):

Differentiate arcsin(2x+1)

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{1-(2x+1)^2} }(2)\]

OpenStudy (anonymous):

@hartnn hey hartnn! can you check my work out?

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{1-(4x^2+4x+1)} }\]

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{-4x^2-4x} }\]

OpenStudy (anonymous):

this is where im stuck, the book has it as

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{-x^2-x} }\]

hartnn (hartnn):

factor out 4 from denominator

hartnn (hartnn):

correct

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

just pull out the four from the radical?

OpenStudy (anonymous):

wouldnt that leave a 2 in the denominator next to the radical/?

hartnn (hartnn):

yes, \(\sqrt{-4x^2-4x}=\sqrt 4(\sqrt{-x^2-x})=2(\sqrt{-x^2-x})\)

hartnn (hartnn):

u have 2 in numerator also!

OpenStudy (anonymous):

OHHHHH!!!!!

OpenStudy (anonymous):

@hartnn your awesome!!!!!!!!!

OpenStudy (anonymous):

:)

hartnn (hartnn):

@Calle87 you too are awesome!